OpenStudy (anonymous):

Which of the following is a linear equation/are linear equations in x1, x2 and x3? A. x1 + 3√x2 + 5x3 = 5 B. √7x1 − 2x2 + √6x3 = √11 1. Only A 2. Only B 3. Both A and B 4. Neither A nor B.

OpenStudy (theviper):

@mukushla plz help ;)

OpenStudy (anonymous):

both of them are linear

OpenStudy (anonymous):

because exponents of \(x_1\) , \(x_2\) and \(x_3\) are 1

OpenStudy (anonymous):

Okay,but the variables in linear equations allowed to have roots and exponents?

OpenStudy (anonymous):

only B.

OpenStudy (anonymous):

you are right, roots are not linear

OpenStudy (anonymous):

@Kudos

OpenStudy (anonymous):

Thanx

OpenStudy (anonymous):

oh sorry that \(\sqrt{x_2}\) so @myko is right

OpenStudy (anonymous):

Agreed only B as linear only has x to power of 1. Taking the second equation B as: sqrt(7) x1 - 2x2 + sqrt(6) x3 = sqrt(11) thus no x variable is contained under any of the sqrt roots!